# Low-Luminosity Type IIP Supernovae from the Zwicky Transient Facility Census of the Local Universe. II: Lightcurve Analysis

### v1

@@ 5.4.2 Lightcurve fitting to radiation-hydrodynamical Type II model grids
5.4.3 Scaling Relations
Scaling relations have been commonly used to infer explosion parameters from lightcurve observables (Popov, 1993; Kasen & Woosley, 2009). Ejecta mass (), progenitor radius (), and explosion energy () are commonly used as the primary independent variables in scaling relations. In this paper, we use scaling relations derived by Goldberg et al. (2019). We use the following relations to constrain the ejected mass () and explosion energy () as a function of progenitor radius (), via the following relations:
@@ 5.4.3 Scaling Relations
The constraints for LLIIP SNe thus obtained are shown in Figure 11.
6 Results
In this section, we provide a summary of the results derived using the analysis techniques discussed in the previous section.
6.1 Observed Parameters
@@
For LLIIP SNe (), the median plateau duration was days, with a range of 64 to 108 days. The median duration of the optically thick phase (OPTd) was days, spanning from 67 to 116 days. The median rise time was days, with a range from 10 to 16 days. The median plateau slope was mag day-1, ranging from to mag day-1. Finally, the change in magnitude between the plateau and the radioactive tail had a median value of mag, with observed values ranging from to mag.
For the full Type IIP SN sample, the median values were days for the plateau duration, days for OPTd, days for the rise time, mag day-1 for the plateau slope, and mag for the plateau-to-tail drop. The steepest drop of mag was seen in SN 2022aagp, shown in Figure 5. A sharp decline from the plateau to the tail is predicted for ECSNe and failed SNe with very low or zero nickel mass (see e.g., Tominaga et al., 2013; Turatto et al., 1998). Other objects with sharp drops include SN 2021tyw ( mag), SN 2022prv ( mag), SN 2023vog ( mag) and SN 2020cxd ( mag).
@@
Brighter SNe tend to have steeper plateau slopes, indicating more rapid declines during the plateau phase. We also note that a major fraction of LLIIP SNe have positive slope. The plateau duration shows only a weak correlation with peak magnitude (). The plateau durations for faint and bright SNe span a broad range and overlap significantly. The physical implications for these results are discussed in Section 7.1.
6.2 Progenitor and explosion parameters
6.2.1 Semi-analytical modeling
For the Nagy & Vinkó (2016) models, the nickel mass is only measured for the group of SNe with a distinct radioactive tail. Based on the semi-analytical modeling, the lowest nickel mass was measured for ZTF20aapchqy/SN 2020cxd ( M⊙). Among the LLIIP sample, the median nickel mass is M⊙, with a range of – M⊙, significantly lower than that of the brighter SNe.
The progenitor radii span – R⊙, with a median of R⊙. The ejecta mass ranges from to M⊙, with a median of M⊙. These explosions have a median kinetic energy of erg and values spanning to erg.
For the full Type IIP SN sample, the corresponding median values are: nickel mass M⊙, radius R⊙, ejecta mass M⊙, and kinetic energy erg. The best-fit values for each SN are listed in Table D. The sample statistics are summarized in Table 3.
@@ 6.2.2 Radiation-hydrodynamical modeling
For the Moriya et al. (2023) models, the ZAMS mass for the LLIIP SNe is typically M⊙, with a median value of M⊙, and a range of – M⊙. The CSM radii span from to cm, with a median of cm. The explosion energy (Eexp) distribution of LLIIP SNe has a median of erg, and values spanning to erg. The typical pre-supernova mass-loss rates inferred have a median of in units, and a range from to .
For the full sample of Type II SNe, the median ZAMS mass is M⊙, with a CSM radius of cm, explosion energy of erg, and mass-loss rate of in units, as summarized in Table 4.
The best-fit lightcurves are shown in Figure 8. After marginalization, we have estimates with confidence intervals (1) for each of these parameters. The distribution of the best-fit parameters and the correlations are shown in Figure 10 and Figure 14. The best-fit values for each SN are listed in Table E. The sample statistics are summarized in Table 4
6.3 Scaling Relations
We investigate the physical properties of the low-luminosity Type IIP (LLIIP) SNe in our sample using scaling relations between plateau luminosity, nickel mass, duration, and progenitor radius. At a representative radius of R⊙, we find that the LLIIP SNe are characterized by relatively low ejecta masses ranging from to M⊙ and low explosion energies between and erg. These results reinforce the picture that LLIIP SNe arise from low-mass progenitors undergoing weak explosions, consistent with those derived from semi-analytical and radiation-hydrodynamical analysis.
7 Discussion
7.1 Observables
7.1.1 LLIIP SNe exhibit less steep gradients, with most displaying upward slopes
@@
7.1.2 The plateau duration and peak magnitude are not correlated
@@
Single-star models predict a strong correlation between peak brightness and plateau duration in Type II SNe (e.g., Kasen & Woosley, 2009). However, the observed lack of such a correlation suggests that the stars undergo substantial pre-supernova mass loss, likely requiring additional processes such as binary interaction to explain the observed diversity, especially at the low-mass end. Eldridge et al. (2018); Hiramatsu et al. (2021); Dessart et al. (2024) also find that binary stripping can explain the duration of plateau seen in Type II SNe lightcurves. This aligns with the increasing observational consensus that a significant number of massive stars are found in binary systems, with many of them undergoing interaction (e.g., Sana et al., 2012). Similarly, Zapartas et al. (2019) estimate that of all Type II SN progenitors undergo binary interaction. Inferring the explosion and progenitor properties from the lightcurves using models that incorporate self-consistent progenitor evolution with binary effects can be used to constrain the binary fraction and extent of stripping among massive stars that explode as core-collapse SNe.
7.2 Correlation of parameters
The dependence of various physical parameters on peak luminosity is shown in Figures 13 and 14. For the Nagy & Vinkó (2016) semi-analytical models, we find a strong correlation between nickel mass and peak absolute magnitude, with a Pearson coefficient of . This follows a log-linear relation:
Kinetic energy also correlates strongly with peak magnitude (), described by:
The strongest correlation is observed between the ratio of kinetic energy to ejecta mass and the peak magnitude ():
These trends show that lower-luminosity SNe tend to have lower nickel masses, lower explosion energies, and lower energy-to-mass ratios.
@@ 7.2 Correlation of parameters
The correlations among physical parameters inferred from the Nagy & Vinkó (2016) models are shown in Figure 15. Kinetic energy and nickel mass are strongly correlated (), following:
Kinetic energy also correlates tightly with ejecta mass (), described by:
Nickel mass and ejecta mass show a weaker positive correlation (). Radius is only weakly correlated with peak luminosity (), and pre-SN mass-loss rates appear uncorrelated.
For the Moriya et al. (2023) radiation-hydrodynamical model fits, significant correlations are also observed. Nickel mass correlates with peak magnitude ():
Explosion energy shows a strong correlation ():
The ratio of explosion energy to ejecta mass exhibits the strongest correlation with peak magnitude ():
Explosion energy correlates with nickel mass ():
ZAMS mass correlates moderately with nickel mass (), radius () and Mr (), but with large scatter. Mass-loss rate () shows no significant correlations ().
7.3 Comparison with literature
A direct comparison with previous studies is challenging due to the heterogeneous nature of the Type IIP SN samples analyzed, as well as the differing methodologies used to infer physical parameters. Below, we briefly summarize relevant literature results for key explosion parameters and compare them with values derived from our semi-analytical modeling and radiation-hydrodynamic modeling.
Rodríguez et al. (2021) studied 110 normal Type II SNe, finding 56Ni masses between – M⊙, with an average of M⊙. Martinez et al. (2022a) reported 56Ni masses of – M⊙ (median M⊙) from 17 objects. Pejcha & Prieto (2015) and Müller et al. (2017) identified similar medians ( M⊙), though their maximum values reached up to M⊙. Meanwhile, Subrayan et al. (2023) measured the nickel mass for 45 SNe, with the lowest nickel mass for one SN being 0.01 and the others ranging between 0.02 and 0.1 . Anderson (2019) measured a broader range of – M⊙, also with a median near M⊙. Hamuy (2003) measured 0.0016 to 0.26 for a sample of 21 Type II SNe. Our semi-analytical modeling yields a high median nickel mass of for the full Type IIP SN sample, spanning 0.0009–0.3098 . The LLIIP subsample shows significantly lower nickel masses, with a median of and a range of 0.0009–0.0420 . The radiation-hydrodynamical fits (Moriya et al., 2023) yield broadly consistent results, with a median nickel mass of for the full sample and for the LLIIP subsample. These results imply that LLIIP SNe are characterized by relatively lower 56Ni yields, in agreement with earlier studies and theoretical predictions for the explosion of low-mass cores (see Section 7.5).
Observationally derived explosion energies typically cluster around – erg (e.g., Martinez et al., 2022a). Martinez et al. (2022a) specifically report a median explosion energy of erg (range – erg). Hamuy (2003) measured erg for a sample of 21 Type II SNe. Utrobin & Chugai (2019) and Pumo et al. (2017) report higher energies, up to – erg. Our semi-analytical modeling yields a high median explosion energy of erg for the full sample, with a range from 0.21–9.94 erg. For the LLIIP subsample, the median energy is substantially lower at erg. We note that the ejecta masses and explosion energies from the semi-analytical fits from Nagy & Vinkó (2016) are higher than those measured in the literature. This is likely due to the two-component density profile assumed in these models. The caveats of the models are further discussed in Section 7.4. The radiation-hydrodynamical fits (Moriya et al., 2023) show relatively lower energies. For the full sample, the median energy is erg, and for the LLIIP events, it is erg, ranging from 0.10–0.28 erg. We note that the lowest energy in the model grid was erg, so it is possible that the explosion energies are even lower.
Typical ejecta masses reported by previous studies, such as Martinez et al. (2022a), are around M⊙ (16–84 percentile range is M⊙). Utrobin & Chugai (2019) report higher ejecta masses, up to M⊙. Hamuy (2003) reported even higher ejecta masses, ranging from . ZAMS masses measured in Subrayan et al. (2023) and Silva-Farfán et al. (2024) ranged from , with most objects at M☉ and a very steep IMF. Our semi-analytical modeling yields a median ejecta mass of for the full sample, with a range from 5.41–24.79 . The LLIIP subsample shows a lower median ejecta mass of , still broadly consistent with the lower end of the literature estimates. This also suggests that the hydrogen envelopes of these massive stars are stripped before they undergo core-collapse. The radiation-hydrodynamical fits (Moriya et al., 2023) give a median ZAMS mass of for the full sample, and for the LLIIP SNe.
Typical progenitor radii reported by previous studies are around – R⊙, with a median of R⊙ (Martinez et al., 2022a). Progenitor radii of 80 to 600 was measured in Hamuy (2003). Irani et al. (2024) find relatively higher radius, with 15 out of 34 SNe II well fit to models with breakout radii cm, while 18 SNe are better fit by a model with a large cm breakout radius. Our semi-analytical modeling yields somewhat smaller radii, with a median of for the full sample and for the LLIIP subsample. These results suggest that LLIIP progenitors may originate from more compact red supergiants or that their outer envelopes have been stripped off via binary interaction at the onset of core-collapse. For comparison, the radius distribution of 74 Galactic RSGs spans a broad range from to R⊙, with a median of R⊙ (Levesque et al., 2005).
Our radiation-hydrodynamical modeling yields median mass-loss rates of for the full sample and for the LLIIP subsample. LLIIP SNe progenitors likely have lower mass-loss prior to explosion. Our results are consistent with the mass-loss rates inferred by Subrayan et al. (2023) and Silva-Farfán et al. (2024), who report values in the range – . The inferred CSM radii have a median value of cm for the full sample and cm for LLIIP events, indicating extended circumstellar environment that could influence early-time lightcurve behavior (e.g., Galbany et al., 2016; Valenti et al., 2016; Förster et al., 2018; Morozova et al., 2018; Bruch et al., 2021, 2023; Jacobson-Galán et al., 2024; Hinds et al., 2025; Jacobson-Galán et al., 2025).
7.4 Caveats of the models
@@
A key challenge arises from the degeneracy between key parameters such as ejecta mass, explosion energy, and progenitor radius. Fitting models to photometry in only a couple of passbands allows multiple combinations of physical parameters to produce similar light curves. This degeneracy limits our ability to extract unique solutions from lightcurve fitting alone as discussed in Singh et al. (2024); Rehemtulla et al. (2025). Goldberg et al. (2019); Goldberg & Bildsten (2020) argue that breaking these degeneracies requires external constraints on at least one physical parameter.
@@
The semi-analytical two-component lightcurve models developed by Nagy & Vinkó (2016) assume spherical symmetry and do not accurately capture the early shock-cooling behavior at days. Accordingly, we restrict our fits to data obtained after 20 days post-explosion. Despite their simplicity, Nagy & Vinkó (2016) and Jäger et al. (2020) find that these models yield physical parameters broadly consistent with more sophisticated hydrodynamical simulations for Type IIP SNe. To test this consistency, Yang et al. (2021) compared fitting results to those from hydrodynamical models such as those in Martinez et al. (2020), and found that, in several cases, the inferred ejecta masses are in reasonable agreement. We find that the semi-analytical fits return higher explosion energies and ejecta mass for most SNe. This could be because of the two-component density profile and a simplified method for computing bolometric lightcurves. Prior works have also noted that lightcurve modeling can systematically overestimate ZAMS masses compared to direct progenitor imaging or nebular spectroscopy (e.g., Utrobin & Chugai, 2008, 2009; Maguire et al., 2010; Sanders et al., 2015).
@@
7.5 What are the progenitors of LLIIP SNe?
7.5.1 7–12 Red Super Giants that undergo core-collapse SNe
We compare our derived explosion parameters to theoretical predictions from neutrino-driven explosion models of core-collapse SNe. Barker et al. (2022) show that the properties of a progenitor’s core can be estimated from optical photometry during the plateau phase alone. A linear correlation between iron core mass and (luminosity 50 days post-explosion) indicates that optical photometry of Type IIP SNe can effectively probe progenitor cores. They find a robust relationship linking progenitor iron core mass to plateau luminosity (see Figure 17). The bolometric luminosities of our LLIIP SN sample align with ZAMS masses below 12 . However, we note that these models assume that there is no CSM interaction. Additionally, using best-fit ZAMS mass estimates from Moriya et al. (2023) mode fits, the ZAMS mass for LLIIP SNe are lower than 11 . Furthermore, ejecta masses inferred from semi-analytical model fitting by Nagy & Vinkó (2016) suggest masses 8.2 . Assuming a neutron star mass of 1.4 , this translates to progenitor masses less than .

### v2

@@ V.4.2 Lightcurve fitting to radiation-hydrodynamical Type II model grids
V.4.3 Scaling Relations
Scaling relations have been commonly used to infer explosion parameters from lightcurve observables (Popov, 1993; Kasen and Woosley, 2009). Ejecta mass (), progenitor radius (), and explosion energy () are commonly used as the primary independent variables in scaling relations. In this paper, we use scaling relations derived by Goldberg et al. (2019). We use the following relations to constrain the ejected mass () and explosion energy () as a function of progenitor radius (), via the following relations:
@@ V.4.3 Scaling Relations
The constraints for LLIIP SNe thus obtained are discussed in Section VI.
VI Results
In this section, we summarize the results obtained using the analysis techniques described above. The best-fit radiation-hydrodynamical values for each SN are listed in Table E, and the sample statistics are summarized in Table 4. Similarly, the best-fit values from the semi-analytical modeling are given in Table D, with the corresponding sample statistics summarized in Table 3.
VI.1 Observed Parameters
@@
For LLIIP SNe (), the median plateau duration was days, with a range of 64 to 108 days. The median duration of the optically thick phase (OPTd) was days, spanning from 67 to 116 days. The median rise time was days, with a range from 10 to 16 days. The median plateau slope was mag day-1, ranging from to mag day-1. Finally, the change in magnitude between the plateau and the radioactive tail had a median value of mag, with observed values ranging from to mag. For the full Type IIP SN sample, the median values were days for the plateau duration, days for OPTd, days for the rise time, mag day-1 for the plateau slope, and mag for the plateau-to-tail drop. The steepest drop of mag was seen in SN 2022aagp, shown in Figure 10. A sharp decline from the plateau to the tail is predicted for ECSNe and failed SNe with very low or zero nickel mass (see e.g., Tominaga et al., 2013; Turatto et al., 1998). Other objects with sharp drops include SN 2021tyw ( mag), SN 2022prv ( mag), SN 2023vog ( mag) and SN 2020cxd ( mag).
@@
Brighter SNe tend to have steeper plateau slopes, indicating more rapid declines during the plateau phase. We also note that a major fraction of LLIIP SNe have positive slope. The plateau duration shows only a weak correlation with peak magnitude (). The plateau durations for faint and bright SNe span a broad range and overlap significantly. The physical implications for these results are discussed in Section VII.
VI.2 Progenitor and explosion parameters
VI.2.1 Nickel mass
For the semi-analytical model fits, the nickel mass is measured only for SNe with a distinct radioactive tail. Within the LLIIP subsample, the median nickel mass is M⊙, spanning 0.0009–0.0420 M⊙. The lowest value in our sample is that of SN 2020cxd, with M⊙. For the full Type IIP sample, the semi-analytical fits yield a median nickel mass of M⊙ (range 0.0009–0.3098 M⊙). The radiation-hydrodynamical fits of Moriya et al. (2023) give broadly consistent values, with a median of M⊙ for the full sample and M⊙ for the LLIIP subsample. Rodríguez et al. (2021) studied 110 events and found 56Ni masses between 0.005–0.177 M⊙ (average M⊙). Martinez et al. (2022b) reported 0.006–0.069 M⊙ (median M⊙) from 17 SNe, while Pejcha and Prieto (2015) and Müller et al. (2017) obtained similar medians (0.03 M⊙) with maxima reaching 0.28 M⊙. Subrayan et al. (2023) measured Ni masses for 45 SNe, with only a single object as low as 0.01 M⊙, while most lie between 0.02 and 0.1 M⊙. Anderson (2019) found a broad distribution (0.001–0.360 M⊙), and Hamuy (2003) reported 0.0016–0.26 M⊙ from 21 SNe. In comparison with these literature samples, the LLIIP SNe in our sample clearly occupy the low end of the nickel-mass distribution. Their very small 56Ni yields are consistent with expectations for explosions of low-mass red supergiant progenitors or events experiencing significant fallback, as discussed in Section VII.5.
VI.2.2 Explosion energy
From the Moriya et al. (2023) radiation-hydrodynamical fits, the LLIIP SNe show low explosion energies, with a median of erg and a range of 0.10–0.28 erg. For the full Type IIP sample, the median energy is erg. In contrast, the semi-analytical Nagy and Vinkó (2016) models yield systematically higher values, with a median of erg for the LLIIP subsample (range 0.21–3.64 erg) and erg for the full sample (range 0.21–9.94 erg). We note that both the ejecta masses and explosion energies from the semi-analytical fits tend to be higher, likely due to the two-component density profile assumed in these models; the caveats are discussed further in Section VII.4. In the literature, Hamuy (2003) measured 0.6–5.5 erg for 21 Type II SNe, while Utrobin and Chugai (2019) and Pumo et al. (2017) obtained up to 1.4–2.0 erg. Martinez et al. (2022b) report a median of 0.61 erg (range 0.30–1.01 erg). The LLIIP SNe in our sample clearly occupy the low-energy end, clustering just above erg, and may in some cases fall below the model grid minimum of erg. This supports interpretations in which LLIIP SNe arise from weak explosions of low-mass cores (see Section VII.5).
VI.2.3 Ejecta mass/ZAMS mass
The semi-analytical models yield ejecta masses of 5.73–12.89 M⊙ for the LLIIP sample, with a median of M⊙. For the full Type IIP sample, the corresponding median ejecta mass is M⊙, spanning 5.41–24.79 M⊙. In earlier studies, Martinez et al. (2022b) report a typical ejecta mass of 9.2 M⊙ (16–84 percentile range 8.2–12.7 M⊙), while Utrobin and Chugai (2019) infer somewhat higher masses up to 14.8 M⊙. Even larger values were found by Hamuy (2003), who estimated ejecta masses in the range 14–56 M⊙ for some Type II SNe. LLIIP SNe lie at the lower end of these distributions, consistent with expectations for low-mass RSG or sAGB progenitors, or for progenitors that have experienced significant envelope stripping prior to core collapse. ZAMS masses inferred in Subrayan et al. (2023) and Silva-Farfán et al. (2024) span 12–16 M⊙, with a steep IMF placing most objects near 12 M⊙. From our radiation-hydrodynamical fits, the LLIIP SNe are inferred to arise from relatively low-mass progenitors, with ZAMS masses typically M⊙ and a median of M⊙. For the full Type II sample, the median ZAMS mass is M⊙. Thus, the LLIIP subsample lies toward the low-mass end of the progenitor-mass distribution, reinforcing the picture that the faintest events originate from the lowest-mass core-collapse progenitors.
VI.2.4 Progenitor radius
The progenitor radii inferred from the semi-analytical fits for the LLIIP SNe span 138–1024 R⊙, with a median of R⊙. For the full Type IIP sample, the corresponding median is R⊙. These radii are smaller than those typically inferred for normal Type II SNe in the literature. For example, Martinez et al. (2022b) report progenitor radii of 460–610 R⊙, with a median of 495 R⊙, while Hamuy (2003) found a broad range of 80–600 R⊙ using hydrodynamical modeling. More recently, Irani et al. (2024) identified a bimodal distribution in breakout radii for 34 SNe II, with roughly half of the sample preferring compact radii ( cm) and the other half favoring larger radii ( cm), consistent with substantial diversity in outer-envelope structure among RSG progenitors. For additional context, the radius distribution of 74 Galactic RSGs spans 100–2135 R⊙ with a median of R⊙ (Levesque et al., 2005). Thus, these results suggest that LLIIP progenitors are either intrinsically more compact RSGs, or that their outer hydrogen envelopes have been partially stripped through binary interaction.
VI.2.5 Progenitor mass-loss
The radiation-hydrodynamical modeling yields median pre-explosion mass-loss rates of M⊙ yr-1 for the full sample, and for the LLIIP subsample. Our results agree well with those of Subrayan et al. (2023) and Silva-Farfán et al. (2024), who report mass-loss rates in the range – M⊙ yr-1 for samples of Type II SNe. The systematically smaller mass-loss rates of the LLIIP subsample may reflect weaker winds or shorter-lived mass-loss episodes prior to explosion. The corresponding CSM radii inferred from the radiation-hydrodynamical fits span – cm, with a median of cm for the full sample and cm for the LLIIP events, indicating an extended circumstellar environment that can influence early-time lightcurve and spectra (e.g., Galbany et al., 2016; Valenti et al., 2016; Förster et al., 2018; Morozova et al., 2018; Bruch et al., 2021, 2023; Jacobson-Galán et al., 2024; Hinds et al., 2025; Jacobson-Galán et al., 2025).
@@ VI.2.5 Progenitor mass-loss
VI.3 Scaling Relations
We investigate the physical properties of the low-luminosity Type IIP (LLIIP) SNe in our sample using scaling relations between plateau luminosity, nickel mass, duration, and progenitor radius. The constraints for LLIIP SNe thus obtained are shown in Figure 11. At a representative radius of R⊙, we find that the LLIIP SNe are characterized by relatively low ejecta masses ranging from to M⊙ and low explosion energies between and erg. These results reinforce the picture that LLIIP SNe arise from low-mass progenitors undergoing weak explosions, consistent with those derived from semi-analytical and radiation-hydrodynamical analysis.
VII Discussion
VII.1 LLIIP SNe exhibit less steep gradients, with most displaying upward slopes
@@
VII.2 The plateau duration and peak magnitude are not correlated
@@
Single-star models predict a strong correlation between peak brightness and plateau duration in Type II SNe (e.g., Kasen and Woosley, 2009). However, the observed lack of such a correlation suggests that the stars undergo substantial pre-supernova mass loss, likely requiring additional processes such as binary interaction to explain the observed diversity, especially at the low-mass end. Eldridge et al. (2018); Hiramatsu et al. (2021a); Dessart et al. (2024) also find that binary stripping can explain the duration of plateau seen in Type II SNe lightcurves. This aligns with the increasing observational consensus that a significant number of massive stars are found in binary systems, with many of them undergoing interaction (e.g., Sana et al., 2012; Duchêne and Kraus, 2013; Kobulnicky et al., 2014; Sana et al., 2014; GRAVITY Collaboration et al., 2018; Bordier et al., 2022; Guo et al., 2022). Also, Zapartas et al. (2019) estimate that – of all Type II SN progenitors undergo binary interaction, based on analytical estimates and population-synthesis simulations that track which progenitors exchange mass with a companion prior to explosion. Inferring the explosion and progenitor properties from the lightcurves using models that incorporate self-consistent progenitor evolution with binary effects will constrain the binary fraction and extent of stripping among massive stars that explode as core-collapse SNe.
VII.3 Correlation of physical parameters
The physical parameters inferred for the sample exhibit a number of clear and physically informative correlations. Figures 13 and 14 summarize the trends between peak luminosity and the key explosion properties. For the Nagy and Vinkó (2016) semi-analytical fits, nickel mass, kinetic energy, and the energy-to-mass ratio all show strong negative correlations with peak absolute magnitude: fainter SNe systematically eject less 56Ni, have lower kinetic energies, and have smaller specific energies. The trends follow the relations:
with Pearson coefficients –0.8. These scalings extend the well-known luminosity–56Ni relation (Hamuy, 2003; Pejcha and Prieto, 2015; Müller et al., 2017; Martinez et al., 2022b) to the faint end and show that LLIIP SNe occupy the faint, low-energy end of a continuous distribution.
@@ VII.3 Correlation of physical parameters
The semi-analytical parameters also correlate with each other. The correlations among physical parameters inferred from the Nagy and Vinkó (2016) models are shown in Figure 15. Nickel mass and kinetic energy are tightly linked (), consistent with the idea that more energetic explosions synthesize more 56Ni. The empirical relation,
is consistent with predictions in neutrino-driven explosion models (Ebinger et al., 2019; Sukhbold et al., 2016; Bruenn et al., 2016, 2023; Sandoval et al., 2021; Bollig et al., 2021; Burrows et al., 2024), as shown in Figure 16. Kinetic energy also correlates strongly with ejecta mass (), following
indicating that explosions with more massive ejecta tend to be more energetic. In contrast, radius and mass-loss rate show only weak or no correlations, implying greater diversity in envelope structure and pre-SN winds.
The correlations inferred from the Moriya et al. (2023) radiation-hydrodynamical models reveal the same trends, but with tighter relationships. Nickel mass and peak magnitude correlate with ,
and explosion energy shows an even stronger dependence on luminosity ():
The specific energy correlates most strongly with luminosity (),
As in the semi-analytical case, explosion energy and nickel mass are tightly linked (),
consistent with neutrino-driven simulations. Mass-loss rate shows no significant correlations (), suggesting that pre-SN winds and envelope stripping vary independently of the core properties that set the SN explosion physics. The overall distribution of the physical parameter correlations is summarized in Figures 13, 14, and the associated correlation matrices.
VII.4 Caveats of the models
@@
A key challenge arises from the degeneracy between key parameters such as ejecta mass, explosion energy, and progenitor radius. Fitting models to photometry in only a couple of passbands allows multiple combinations of physical parameters to produce similar light curves. This degeneracy limits our ability to extract unique solutions from lightcurve fitting alone as discussed in Singh et al. (2024); Rehemtulla et al. (2025). Goldberg et al. (2019); Goldberg and Bildsten (2020) argue that breaking these degeneracies requires external constraints on at least one physical parameter.
@@
The semi-analytical two-component lightcurve models developed by Nagy and Vinkó (2016) assume spherical symmetry and do not accurately capture the early shock-cooling behavior at days. Accordingly, we restrict our fits to data obtained after 20 days post-explosion. Despite their simplicity, Nagy and Vinkó (2016) and Jäger et al. (2020) find that these models yield physical parameters broadly consistent with more sophisticated hydrodynamical simulations for Type IIP SNe. To test this consistency, Yang et al. (2021) compared fitting results to those from hydrodynamical models such as those in Martinez et al. (2020), and found that, in several cases, the inferred ejecta masses are in reasonable agreement. We find that the semi-analytical fits return higher explosion energies and ejecta mass for most SNe. This could be because of the two-component density profile and a simplified method for computing bolometric lightcurves. Prior works have also noted that lightcurve modeling can systematically overestimate ZAMS masses compared to direct progenitor imaging or nebular spectroscopy (e.g., Utrobin and Chugai, 2008, 2009; Maguire et al., 2010; Sanders et al., 2015).
@@
VII.5 What are the progenitors of LLIIP SNe?
VII.5.1 7–12 Red Super Giants that undergo core-collapse SNe
We compare our derived explosion parameters to theoretical predictions from neutrino-driven explosion models of core-collapse SNe. Barker et al. (2022) show that the properties of a progenitor’s core can be estimated from optical photometry during the plateau phase alone. A linear correlation between iron core mass and (luminosity 50 days post-explosion) indicates that optical photometry of Type IIP SNe can effectively probe progenitor cores. They find a robust relationship linking progenitor iron core mass to plateau luminosity (see Figure 17). The bolometric luminosities of our LLIIP SN sample align with ZAMS masses below 12 . However, we note that these models assume that there is no CSM interaction. Additionally, using best-fit ZAMS mass estimates from Moriya et al. (2023) mode fits, the ZAMS mass for LLIIP SNe are lower than 11 . Furthermore, ejecta masses inferred from semi-analytical model fitting by Nagy and Vinkó (2016) suggest masses 8.2 . Assuming a neutron star mass of 1.4 , this translates to progenitor masses less than .